# Two Minds Better Than One: Collaborative Reward Modeling for LLM Alignment

### v1

@@ 5 Related Work
Despite the substantial effort of Reinforcement Learning from Human Feedback (RLHF) Ouyang:2022:RLHF , realizing an ideal proxy RM is challenging. Existing RM face issues like reward misgeneralization di:2022:goal_mis ; gao:2023:ScalingLawsReward ; Qiu:2024:RewardTopological ; chen:2024:DiscriminRM ; bukharin:2025:advRM and reward hacking Skalse:2022:Hack ; Miao:2024:InfoRM ; Fu:2025:Hack . These challenges motivate the alternatives of RLHF, which can be broadly categorized into two main approaches: RLHF with explicit reward models and with implicit reward models. The first category focuses on calibrating the reward signal by designing regularization techniques. These include mitigating length bias Chen:2024:ODIN and confidence bias Leng:2024:Confidence inherent in reward models, regularizing score based on internal state states zhang:2024:OvercomingRewardOveroptimization ; shen:2024:ImprovingReinforcementLearning ; miao:2025:EnergyLossPhenomenon , ensembling multiple reward models coste:2023:REWARDMODELENSEMBLES ; eisenstein:2024:HelpingHerdingReward ; yan:2024:RewardRobustRLHFLLMs ; Wang:2024:secretsII , and scaling up dataset and parameters gao:2023:ScalingLawsReward ; zhai:2023:UncertaintyPenalizedReinforcementLearning ; Wang:2024:MoeRM ; Liu:2024:RRM . The second directly optimizes the policy LLM on preference datasets without explicit reward model. Representative methods like DPO Rafailov:2023:DPO derive the closed-form solution implied by the Bradley-Terry model. Recently, several implicit reward model methods emerged, including IPO Azar:2024:IPO , KTO Ethayarajh:2024:KTO , ORPO Hong:2024:ORPO , and SimPO Meng:2024:Simpo . However, methods with implicit reward suffer from sub-optimality Xu:2024:dpo_vs_ppo ; ivison:2024:DPO_vs_PPO and distribution shift Guo:2024:dpo_ood ; Yang:2025:DPO_Shift .
Robust Preference Alignment. Recent research have increasingly recognized the detrimental effects of NPs gao:2024:ImpactPreferenceNoise ; Wang:2024:secretsII ; IM:2025:NP_Theory , elicting a series of robust preference alignment methods Mitchell:2023:cDPO ; Chowdhury:2024:rDPO ; Alexander:2024:Corrupted_Human ; Liang:2024:ROPO ; Wu:2024:DrDPO . IM:2025:NP_Theory theoretically derives an upper bound on the generalization error under certain noisy ratio, and analyzes how generalization error changes as the noise increases. To mitigate the negative effect of NPs on policy models, cDPO Mitchell:2023:cDPO designs a weighted binary cross-entropy loss by incorporating a conservative target distribution to account for NPs. Dr.DPO Wu:2024:DrDPO introduces an additional parameter $\beta^{\prime}$ to balance the importance of preference pairs. Furthermore, ROPO Liang:2024:ROPO designs a conservative gradient weighting strategy to suppress the influence of NPs. Recent work Wang:2024:secretsII has examined noisy preferences in open-source datasets, proposes to filter noisy preferences based on the consistency between reward models. In contrast, our work is the first to systematically analyze the intrinsic characteristics of preference pairs and perform online noise filtering based on these characteristics.
@@ 6 Conclusion, Limitation and Future Work
Limitation and Future Work. Our work offers a promising solution to mitigate noisy preferences in the optimization of reward models, though there remains room for improvement. Firstly, due to limited computational resources, we implement experiments using LLaMa3-3B as the base model. Although extensive experiments validate the proposed method, future work is suggested to enhance effectiveness by scaling the parameters of the language model. Secondly, we aim to generalize our framework to other established alignment methods, such as SimPO Meng:2024:Simpo , and extend the scope of our alignment framework to other domains. Additionally, it is worth exploring the introduction of game theory Neumann:2007:Game_Theory ; Shoham:2008:multiagent into reward modeling.
@@ B.1 Tasks and Dataset
8:         Model $r_{\phi}$ peer review: $\mathcal{B}_{\phi}=\operatorname{argmax}_{\mathcal{B}_{\psi}\subseteq\mathcal{B},|\mathcal{B}_{\psi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\phi};\mathcal{B}\right),$
9:         Model $r_{\psi}$ peer review: $\mathcal{B}_{\psi}=\operatorname{argmax}_{\mathcal{B}_{\phi}\subseteq\mathcal{B},|\mathcal{B}_{\phi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\psi};\mathcal{B}\right),$
@@
B.2 Implementation Details
@@ B.3 Baselines
Settings Opponent Arena-Hard MT-Bench
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
0% Filpped DPO 34%38%28% 45%29%26%
cDPO 58%30%12% 68%21%11%
rDPO 33%39%28% 51%23%26%
ROPO 32%42%26% 48%26%26%
20% Filpped DPO 38%41%21% 46%29%25%
cDPO 61%28%11% 72%18%10%
rDPO 32%45%23% 34%36%30%
ROPO 42%39%19% 48%25%27%
40% Filpped DPO 51%35%14% 61%30%9%
cDPO 72%22%6% 47%41%12%
rDPO 46%35%19% 59%11%30%
ROPO 53%32%15% 50%35%15%
Table 10: Ablation analysis of two policy LLMs within CRM.
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
0% Filpped DPO 80%12%8% 65%12%23% 51%35%14%
CRM-Model 1 84%8%8% 73%11%16% 52%31%17%
CRM-Model 2 80%14%6% 69%14%17% 48%33%19%
20% Filpped DPO 77%13%10% 63%13%24% 40%32%28%
CRM-Model 1 82%11%7% 73%10%17% 48%33%18%
CRM-Model 2 85%10%5% 70%11%19% 50%34%16%
40% Filpped DPO 70%17%13% 61%30%9% 29%35%36%
CRM-Model 1 80%9%11% 71%12%17% 43%36%21%
CRM-Model 2 81%12%7% 73%13%14% 47%32%21%

### v2

@@ 5 Related Work
Despite the substantial effort of Reinforcement Learning from Human Feedback (RLHF) [2], realizing an ideal proxy RM is challenging. Existing RM face issues like reward misgeneralization [16; 17; 19; 18; 39] and reward hacking [40; 25; 41]. These challenges motivate the alternatives of RLHF, which can be broadly categorized into two main approaches: RLHF with explicit reward models and with implicit reward models. The first category focuses on calibrating the reward signal by designing regularization techniques. These include mitigating length bias [42] and confidence bias [43] inherent in reward models, regularizing score based on internal state states [44; 15; 45], ensembling multiple reward models [46; 47; 48; 9], and scaling up dataset and parameters [17; 49; 50; 51]. The second directly optimizes the policy LLM on preference datasets without explicit reward model. Representative methods like DPO [26] derive the closed-form solution implied by the Bradley-Terry model. Recently, several implicit reward model methods emerged, including IPO [52], KTO [53], ORPO [54], and SimPO [55]. However, methods with implicit reward suffer from sub-optimality [56; 57] and distribution shift [58; 59].
Robust Preference Alignment. Recent research have increasingly recognized the detrimental effects of NPs [14; 9; 60], elicting a series of robust preference alignment methods [20; 21; 61; 23; 22]. [60] theoretically derives an upper bound on the generalization error under certain noisy ratio, and analyzes how generalization error changes as the noise increases. To mitigate the negative effect of NPs on policy models, cDPO [20] designs a weighted binary cross-entropy loss by incorporating a conservative target distribution to account for NPs. Dr.DPO [22] introduces an additional parameter $\beta^{\prime}$ to balance the importance of preference pairs. Furthermore, ROPO [23] designs a conservative gradient weighting strategy to suppress the influence of NPs. Recent work [9] has examined noisy preferences in open-source datasets, proposes to filter noisy preferences based on the consistency between reward models. In contrast, our work is the first to systematically analyze the intrinsic characteristics of preference pairs and perform online noise filtering based on these characteristics.
@@ 6 Conclusion, Limitation and Future Work
Limitation and Future Work. Our work offers a promising solution to mitigate noisy preferences in the optimization of reward models, though there remains room for improvement. Firstly, due to limited computational resources, we implement experiments using LLaMa3-3B as the base model. Although extensive experiments validate the proposed method, future work is suggested to enhance effectiveness by scaling the parameters of the language model. Secondly, we aim to generalize our framework to other established alignment methods, such as SimPO [55], and extend the scope of our alignment framework to other domains. Additionally, it is worth exploring the introduction of game theory [62; 63] into reward modeling.
@@ B.1 Tasks and Dataset
8:         Model $r_{\phi}$ peer review: $\mathcal{B}_{\phi}=\operatorname{argmax}_{\mathcal{B}_{\phi}\subseteq\mathcal{B},|\mathcal{B}_{\phi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\phi};\mathcal{B}\right),$
9:         Model $r_{\psi}$ peer review: $\mathcal{B}_{\psi}=\operatorname{argmax}_{\mathcal{B}_{\psi}\subseteq\mathcal{B},|\mathcal{B}_{\psi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\psi};\mathcal{B}\right),$
@@ B.2 Implementation Details
Settings Opponent Arena-Hard MT-Bench
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
0% Filpped DPO 34%38%28% 45%29%26%
cDPO 58%30%12% 68%21%11%
rDPO 33%39%28% 51%23%26%
ROPO 32%42%26% 48%26%26%
20% Filpped DPO 38%41%21% 46%29%25%
cDPO 61%28%11% 72%18%10%
rDPO 32%45%23% 34%36%30%
ROPO 42%39%19% 48%25%27%
40% Filpped DPO 51%35%14% 61%30%9%
cDPO 72%22%6% 47%41%12%
rDPO 46%35%19% 59%11%30%
ROPO 53%32%15% 50%35%15%
Table 6: Ablation analysis of two policy LLMs within CRM.
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
0% Filpped DPO 80%12%8% 65%12%23% 51%35%14%
CRM-Model 1 84%8%8% 73%11%16% 52%31%17%
CRM-Model 2 80%14%6% 69%14%17% 48%33%19%
20% Filpped DPO 77%13%10% 63%13%24% 40%32%28%
CRM-Model 1 82%11%7% 73%10%17% 48%33%18%
CRM-Model 2 85%10%5% 70%11%19% 50%34%16%
40% Filpped DPO 70%17%13% 61%30%9% 29%35%36%
CRM-Model 1 80%9%11% 71%12%17% 43%36%21%
CRM-Model 2 81%12%7% 73%13%14% 47%32%21%